OpenStudy (anonymous):

a^2-10 factor and tell me how it works, please!

OpenStudy (anonymous):

Move to right hand side \[a^2=10\] Root each side \[a=\pm\sqrt{10}\] Because \(a=\pm\sqrt{10}\), then the factors must be \[(a+\sqrt{10})(a-\sqrt{10})\]